Find 11, 12 and V1 in the following circuit: 422 613 32
Answer #1

her — ar — Problem Adding Series Connected resistors. - G+6=12 25 ur are Connected in Series 32 & 30 264262 figure at bor arfoltage across 32 resistor = 232*3 = +2+3=3-64 Apply kve to the loop shown in figure, 1 tv, this HGR =0. V = -(3.6+3.6) = -7-2